Ideally, when a thermometer is used to measure the temperature of an object, the temperature of the object itself should not change. However, if a significant amount of heat flows from the object to the thermometer, the temperature will change. A thermometer has a mass of 29.2 g,.a specific heat capacity of c = 881 J/(kg Се), and a temperature of 12.1 °C. It is immersed in 138 g of water, and the final temperature of the water and thermometer is 62.4 °C. What was the temperature of the water in degrees Celsius before the insertion of the thermometer?
Answer #1

thermometer specific heat capacity of C= 881J/(kg C°)

H2O a specific heat capacity of C = 4186 J/(kg C°)

heat gained = heat lost = Q = mCθ°

where θ° is the temperature difference, m = mass in kg

=> heat gain = 0.0292(881)[62.4 - 14.4]° = heat loss = 0.138(4186)[T - 62.4]

=> 1234.8096 = 577.668T - 36046.4832

=> 577.668T = 37,281.2928

or T ~= 64.537°C


hope this helps